OpenStudy (anonymous):

if f(x) is a linear function such that f(2)=5 and f(4)=13 then f(x)=?

OpenStudy (anonymous):

@amriju reply plz

OpenStudy (amriju):

give me a minute to try...

OpenStudy (anonymous):

F(x) is a linear function, meaning that f(x)=ax+b. F(2)=5. Meaning that F(2)=2a+b and 2a+b=5. F(4)=13, meaning that F(4)=4a+b and 4a+b=13. Now we have two equations to solve simultaneously. 2a+b=5 4a+b=13 Minus 2 from 1. Leaving: -2a=-8 a=4 2(4)+b=5 8+b=5 b=-3 Now we know that that F(x)=4x-3

OpenStudy (amriju):

how cud u be sure of a linear func.. @Henry.Lister

OpenStudy (amriju):

cud be any func

OpenStudy (anonymous):

Read the question.

OpenStudy (amriju):

:P

OpenStudy (amriju):

very sorry........

OpenStudy (anonymous):

No problem.